OpenStudy (anonymous):

Determine whether ΔABC ~ ΔFED. If they are similar, what postulate or theorem proves their similarity? Choices : Side-Angle-Side Similarity Postulate Side-Side-Side Similarity Postulate Similarity cannot be determined. Angle-Angle Similarity Postulate

OpenStudy (anonymous):

Side-side-side?

OpenStudy (eyust707):

Yes! I believe so.

OpenStudy (eyust707):



OpenStudy (anonymous):

yes, because you only refer to the sides

OpenStudy (anonymous):

Eyust, \[\frac{45}{20} = \frac{36}{16} =? \frac{27}{12}\]

OpenStudy (eyust707):